OpenStudy (anonymous):

Solve: 9 - 30x + 25 = 0 a. {5/3} b. {-5/3} c. {-5/3, 5/3} d. {-3/5, 3/5}

OpenStudy (anonymous):

Is the question correct? 'Cause if it is, my answer is none of the above.

OpenStudy (amistre64):

prolly 9x^2

OpenStudy (anonymous):

do you mean 9x^2?

OpenStudy (anonymous):

Yeah it's 9x^2 ...Sorry guys. :[

OpenStudy (amistre64):

quad formula is the most plug and play outcome

OpenStudy (anonymous):

(3x-5)(3x-5)=0

OpenStudy (anonymous):

A. :)

OpenStudy (amistre64):

9x^2 - 30x + 25 \[\frac{30}{2(9)}\pm\frac{\sqrt{30^2-4(9)(25)}}{2(9)}\]

OpenStudy (anonymous):

just take one 3x-5=0 3x=5 x=5/3